help

A private opinion poll is conducted for a politician to determine what proportion of the population favors decriminalizing marijuana possession. How large a sample isneeded in order to be 98% confident that the sample proportion will not differ from the true proportion by more than 5%?
Answer #1

uploaded image

α=0.02

tα/2s/√n=0.05xbar

n=(tα/2s/0.05xbar)2

answered by: odie
Answer #2
a=0.02, |Z(0.01)|=2.33 (check standard normal table)

SO n=(Z/E)^2*p*(1-p)
=(2.33/0.05)^2*0.5*0.5
=542.89

Take n=543